OpenStudy (anonymous):

Solve the following system of equations 2x - 3y + z = -1 2x - 2y - 2z = -8 x + 5y - 3z = 2

OpenStudy (anonymous):

Would it be (-1,2,8)?

hero (hero):

Did you actually solve it to get that or are you just guessing?

OpenStudy (anonymous):

guessing lol. I'm doing my exams and it's timed and I'm getting low on time here

hero (hero):

No, you can't guess. Just use wolfram alpha

OpenStudy (anonymous):

Guessing would work, but need to make some smart guess to though.

OpenStudy (anonymous):

If he has an exam, he has to understand and wolfram alpha doesn't help too much.

hero (hero):

x = 1 y = 2 z = 3

hero (hero):

You can use wolfram alpha to understand. You just have to know how to use it

OpenStudy (anonymous):

Wolfram would help you see if you are doing it right or not. It is a good tool to assist you while you are learning system of equation concepts.

OpenStudy (anonymous):

2x = 3y - z - 1 2x = 2y + 2z - 8 2x = -10y + 6z + 4 (multiply by 2 for the whole eqn) So, 3y - z - 1 = 2y + 2z - 8 = -10y + 6z + 4 go on doing the simultaneos eqn by any method u like and i guess u'll get the answer right. Pls do tell me if there's a btr method, I'm juz finished Form 5 (or O lvl for other countries) and thts what i guessed i've learned to solve this kind of Q.